OpenStudy (anonymous):

how can I solve this equation e^(x-2)= e^(4+x^1/2)

OpenStudy (anonymous):

equate the exponents

OpenStudy (anonymous):

its a quadratic eqn , if you set u=x^(1/2)

OpenStudy (anonymous):

u^2 -2 = 4 +u

OpenStudy (anonymous):

etc, you can do the rest

OpenStudy (anonymous):

and you'll find one possible result which is : 9

OpenStudy (anonymous):

Thx! for the help